# Towards AI-Architecture Liberty: A Comprehensive Survey on Designing and Collaborating Virtual Architecture by Deep Learning in the Metaverse

### v3

@@ 3.3.2. Architectural Information Extraction from Latent Space
For research on architectural generation, VAEs extract information through the latent space with a pair of encoder-decoders, which is compatible with various design purposes but is seldom utilized in architectural design due to high technical access. As one pioneering research of 3D generation, Miguel et al. used labeled connectivity vectors extracted from rectangular cube structure ”3D-canvas” as data representation because of the usability and compact regarding 3D feature structure (Miguel et al., 2019). The outcome is achieved by 3D voxelized wireframes of multiple architectural forms with different styles and strengths through learning continuous latent distributions of VAE. Based on this work, Miguel et al. further tested parametric augmentation for a larger dataset of 3D geometries (Miguel et al., 2020). By contrast, some work concentrates on post-integration assisting training according to the designer’s preference after VAE generation in the whole production. For instance, Kahraman et al. produced a series of chairs with 3D voxelization while comparing multi-object VAE and GAN (Kahraman et al., 2021). This application discussed pre-defined criteria ranging from leisure to work scenarios.
3.4. 3D-Aware Image Synthesis
@@
Most 3D-aware image synthesis methods utilize GAN-based models to sample latent vectors and decode them into target 3D representations. Thus, it offers 3D view-consistent rendering, efficient presentation, and interactive editability, as well as integration capabilities with other deep learning generation techniques (Shi et al., 2023b), which could contribute to virtual architectural generation.
Method Names Publication & Year 3D Repre- sentations Single/multiple-view Geometry Editability
Controllability
Highlight
Camera
Object
NeRF  (Gu et al., 2021) CVPR 2019 Neural field
multiple
✓relighting
3D Novel View Synthesis
HoloGAN  (Nguyen-Phuoc et al., 2019) CVPR 2019 Voxel grid
single
Pi-GAN  (Chan et al., 2021) CVPR 2021 Neural field
single
position
Giraffe  (Niemeyer and Geiger, 2021) CVPR 2021 Neural field
single
StyleSDF  (OrEl et al., 2022) CVPR 2022 Neural field
single
StyleNeRF  (Gu et al., 2021) arXiv 2021 Neural field
single
DreamField  (Jain et al., 2022) CVPR 2022 Neural field
multiple
CLIP: Text input
DreamFusion  (Poole et al., 2022) arXiv 2022 Neural field
single
Text input
CLIP-NeRF  (Wang et al., 2022) CVPR 2022 Neural field
single
CLIP: Text input
Table 5. An overview of 3D generative approaches of 3D-Aware Image Synthesis. Single/multiple represents the result generated by a single image adopting a sample of single-view or multiples image adopting multiple-view images. Geometry indicates whether this method allows the export to mesh. Editability indicates whether this generation process enables editing, such as composing objects in the scene. 3D-aware image synthesis is performed by controllability, including camera pose, position or object pose, location, relighting, and so on.
3.4.1. 3D-Aware Image Synthesis and Editability
@@
3D-aware synthesis shows promise for virtual architecture regarding representation conversion, controllability, and multi-modality with linguistic descriptions. As OrEl et al. demonstrated in StyleSDF, converting implicit representations to explicit geometry enables designers to comprehend and edit architectural elements with precision visually (OrEl et al., 2022). Similarly, Gu et al. showcased in StyleNeRF how explicit control over style attributes allows for creating diverse architectural designs with semantic editing capabilities (Gu et al., 2021). In recent advancements, Meng et al. have introduced a configurative Colab with user-friendly interaction for creating buildings based on DreamField (Jain et al., 2022), supporting conditional text input, parameter adjustment, and style attributes  [^16] Source:https://github.com/shengyu-meng/dreamfields-3D. This implementation underscores the practical application of 3D-aware synthesis in architectural design, providing designers with intuitive tools to realize their creative vision. However, concerns arise regarding effectiveness limited to simple objects due to non-design datasets. Specifically, this method has limitations in achieving high-resolution and precise internal structures for generating buildings. Additionally, efficiency is significantly reduced when generating complex architectural structures with both internal structures and appearances due to the characteristics of deep generative models.
3.5. Emerging Generation Based on Diffusion Model
@@
3D diffusion is capable of high quality with fine details in a robust and controlled way for 3D generation, particularly complex shapes with intricate details and specific attributes. Thus, its controllability and descriptive input will contribute to deep-learning-assisted virtual architectural generation in the future.
3.5.1. 3D Diffusion
Recently, 3D diffusion has gained a growing interest in generating 3D shapes since it outperforms GANs in fidelity regarding intricate details and sharp edges, as well as stability during training. This superiority arises from its fine-grained control over the generation process, enabling specific attribute control or smooth interpolation between shapes.
Despite the flexibility of conditional diffusion sampling, traditional diffusion models, as a form of deep generative model (DGM), are limited to sampling pixels. In contrast, Ben et al. have innovatively applied diffusion models to denoise images for high-quality 3D-aware image synthesis, as demonstrated by DreamFusion (Poole et al., 2022). DreamFusion bypasses the need to process large amounts of 2D image data by directly generating 3D models. Based on this work, Magic3D has addressed efficiency issues related to low-resolution geometry and textures, as well as computational and memory-intensive operations, surpassing DreamFusion in performance (Lin et al., 2023). Another work named LION demonstrates higher quality performance by leveraging diffusion models in combination with a VAE, offering greater flexibility in operation and application due to conditional synthesis and shape interpolation (Zeng et al., 2022). Unlike most existing diffusion-based probabilistic models (DDPMs), PVD employs a unified probabilistic formula to generate high-fidelity 3D shapes from a single-view depth scan of real objects (Zhou et al., 2021).
@@
Diffusion models facilitate the generation of 3D shapes with outperformed controllability in two ways: modifying style attributes and inputting conditioned descriptive text. Firstly, most methods allow modification of specific style attributes, such as shape and texture, during the generation process (Gu et al., 2021; Niemeyer and Geiger, 2021; Jain et al., 2022). This progress suggests that diffusion models offer a robust and controlled approach to 3D shape generation, particularly for complex shapes with intricate details and specific attributes. Secondly, some burgeoning methods feature text-to-3D generation with human interpretative intention by inputting description (Chen et al., 2019; Poole et al., 2022; Liu et al., 2022; Lin et al., 2023; Nichol et al., 2022b). The earliest research we tracked is text2shape (Chen et al., 2019), in which 3D models with color and shape paired with natural language formed datasets to build implicit semantic links. Recent research has made a remarkable breakthrough in associating text and 3D models with unsupervised learning. Similar to 3D-aware synthesis, most methods utilize CLIP with unsupervised learning (Lin et al., 2023; Nichol et al., 2022b).
Moreover, this approach broadens the usage case for other layman users by interpreting textual prompts. For instance, Magic 3D has developed a toolkit that offers advanced control over image instances of 3D generation regarding stylization and content based on text prompts (Lin et al., 2023). Therefore, the text-to-3D approach democratizes 3D geometry generation, providing access for individuals with varying levels of expertise to produce creatively.
3.5.2. 3D Form Driven by Text in Architecture
Despite limited applications in 3D form generation, this emerging diffusion model is one of the most promising and efficient models in design application due to text-based description. Some initial research in design has leveraged language regarding human interpretation, interaction usability, and integrating multimodality. An earlier work is that Zhang developed a framework capable of compiling analytic semantic information of the input texts, then generating 3D architectural form according to descriptive language (Zhang, 2020). In this framework, different usages of spaces have been trained with adjacent matrices to understand the linguistic instructions. Although little research utilizes this model, pioneering design works utilizing the diffusion model have proliferated recently. For instance, AIG studio’s work  [^17] Source: https://www.bilibili.com/video/BV1Qb411Z7UP/ has achieved advanced editing to train urban morphology heatmaps from 3D models utilizing Lora models through stable diffusion. The resultant grayscale images, embodying height data, facilitate seamless mesh model conversion within the same modeling framework. Comparably, the latest work mainly contributes to integrating tools to improve common tedious workflows, such as constant transforming models between multiple software or methods. In this work, Guida developed a user interface integrating 3D form generation for designers to embrace multi-modality (Guida, 2023).
4. Characteristics of Generated Virtual Architecture in Designer-Machine Learning Collaboration
@@ 4.2. Scope
The degree of scope depends on two aspects: the involvement of process scope in the design and deep learning-assisted design or human-assisted design. Firstly, involvement in more design phases represents a wider scope, including data training and processing, solution set generation, and evaluation (as illustrated in Section 2.3). For instance, Mueller conducted the entire scope of design work involving pre-processing datasets, refining and fine-tuning models, and evaluation based on GAN to test the impact of deep learning on design output (Mueller, 2023). By comparison, Zhang et al. merely applied GANs to generate a set of section images for 3D transposition without involving other phases, representing a specific scope (Zhang, 2019).
Secondly, the work involving human-assisted design has a wider scope than deep learning-assisted design regarding design generation since it accommodates additional processes beyond design generation, such as dataset augmentation (Miguel et al., 2019) and semantic labeling (Zhang, 2020). For instance, Huang et al. extended architects’ capabilities by combining GAN generation with analytical interpretation processes of NLP (natural language processing) instead of solely focusing on design generation (Huang et al., 2021). Such research addresses the limitations of DL-assisted design generation and compensates for them by incorporating a wider scope of design perspectives. In contrast, most studies relying on generative images only investigate DL-assisted design without addressing other concerns (Zhang, 2019; Yu, 2020).
@@ 4.3. Operability
Operability refers to the extent of technical or design operation involved in design collaboration with DGMs. Low operability indicates the straightforward application of DGMs to design tasks, while high operability signifies the ability to operate in a work within the technical space with improved functionality and efficiency. Technical operations encompass various types across different models, including fine-tuning model parameters (Mueller, 2023), refining training procedures (Ondrej Vesel‘y, 2022; Çakmak, 2022; Mueller, 2023; Sebestyen et al., 2021), customizing architecture (Kahraman et al., 2021), and adapting methodology applicability (Huang et al., 2023).
The degree of operability is influenced by the complexity of technical operations and the applicability of methodology. Firstly, the different focuses of design work contribute to varying levels of technical operations and the applicability of the methodology. For instance, Veselý demonstrated high operability by unifying methodology applicability and fine-tuning GAN models for different scales and scopes, including adjusting hyperparameters and optimizing the training process to achieve better convergence and output quality (Ondrej Vesel‘y, 2022). Similarly, Huang et al. explored the latent space potential by connecting structured segmentation of text and visual outputs in generative images of the diffusion model, establishing robust analytical frameworks operation (Huang et al., 2023). Secondly, different designs exhibit varying levels of applicability due to model divergence. Generally, VAE requires relatively high operability from professionals since it processes low-dimensional 3D feature data of buildings. In contrast, diffusion and GAN offer various levels of operability with flexibility ranging from application to fine-tuning, while 3D-aware synthesis requires an intermediate level of operability.
@@ 4.4. Autonomy
Autonomy refers to the level of automation in the generation process facilitated by deep learning techniques. High autonomy indicates a largely automated process driven by a unified workflow, while low autonomy involves manual intervention, requiring the connection of workflows and greater ease of use. In architectural design, most works exhibit relatively low autonomy to ensure quality and design accuracy under the guidance of designers. For example, Mueller validated the methodology across diverse scenarios by selectively retrieving spatial data and comparing the performance of proposed generative models (Mueller, 2023). However, recent advancements show a proliferation of high autonomy, especially in 3D generation, focusing on task efficiency (Kim et al., 2020; Lab, 2023). Nevertheless, there’s a growing trend towards higher autonomy in design work to broaden access and enhance usability. For instance, Guida developed an integrated tool within the visual scripts of Grasshopper with high autonomy, combining all processes and modules in architectural design, providing convenient input and output modules across different approaches (Guida, 2023).
@@ 5. Research Agendas
In the previous sections, we delve into the urgent need for HCI and DL-assisted methods in generated virtual architecture. Through thoroughly examining the literature, we elucidate four key focuses, as mentioned in Section 2.3. These key focuses serve as the foundation for our research agendas, formulated from a perspective of high-level questions (Fig. 9). Our investigation of current literature and works also reflects on these questions. Therefore, to promote research in this area, we advocate the following research agendas.
Figure 9. A summarizing diagram covering the connection between key focuses (Section 2), DGMs (Section 3), and characteristics of generation approaches (Section 4). Accordingly, high-level issues lie behind four key focuses in our survey form further research agendas (Section 5).
@@ 5.1. Agency Between Designer and Machine in Data Processing and Training
Agency between humans and machines emerges as a significant feature in AI-assisted design and human-AI collaboration. On the one hand, a human-driven agency faces challenges due to low-resolution and incomplete datasets, requiring human intervention (Shi et al., 2023a). On the other hand, an AI-driven agency raises concerns about intellectual property issues but offering higher efficiency (Guida, 2023). Our analysis of the survey identifies two main barriers to achieving a balance between human and AI agencies: flexible datasets and human-assisted machine learning.
@@ 5.2. Augmenting Communication Between Design and Artificial Intelligence
To benefit the design project’s fitting purpose, the HCI researcher and designer should augment communication with each other by promoting tools or enhancing workflows between design applications and deep learning. We found two aspects closely related to this agenda: explainability in human-AI collaboration and editability in the generative model.
Explainability. Explainability is one recognized principle of communication within human-AI collaboration, essential for fostering trust in AI systems and guiding the design process and outcomes (Mohseni et al., 2021). However, a significant disparity exists in the developmental progress between HCI and the design field. For instance, HCI researchers have made tremendous strides in responsible AI for explainability, focusing on user experience, trustworthiness, and collaboration (Saeed and Omlin, 2023). The design field lags due to existing challenges and a lack of expertise among designers in negotiating input, process, and controlling output. According to our analysis of the survey, no one has addressed this critical issue. Regardless, recent studies (N=2) have primarily investigated methods for controlling output by manipulating detail in algorithms within generative models to achieve greater controllability in design outputs. Therefore, future research in HCI and design should delve deeper into explainability in design and generation processes building upon design principles.
Editability. Editability is the other key to communication, relying on timely feedback and adjustments to model 3D objects. Similarly, there is a big gap in development progress between the generation technique and design fields. Prior work has demonstrated editing autonomy in 3D shape generation techniques available to users. For example, Liu et al. proposed a user-centric 3D shape generation method by drawing a target model as 3D voxel grids (Liu et al., 2017). While a few works demonstrate some indirect approaches to editing 3D models with implicit representations (Hao et al., 2020; Ibing et al., 2021; Deng et al., 2021b; Zheng et al., 2021). According to our survey, although none of the architectural works mainly contribute to editing in 3D generation, one indicated a potential for variability in 3D form. This work generates infinite continuous forms by interpolating 3D features among multiple wireframe representations with compact implicit functions (Miguel et al., 2019). Through this interpolation with implicit representations, we could envision a broad use for virtual architecture in supporting itself with real-time dynamic morphology. Thus, the editability potential of implicit representation incorporating building form launches the future agenda for deep learning virtual architecture design.

### v4

@@ 3.3.2. VAEs Utilized in Architecture
For research on architectural generation, VAEs extract information through the latent space with a pair of encoder-decoders, which is compatible with various design purposes but is seldom utilized in architectural design due to high technical access. As one pioneering research of 3D generation, Miguel et al. used labeled connectivity vectors extracted from rectangular cube structure ”3D-canvas” as data representation because of the usability and compact regarding 3D feature structure (Miguel et al., 2019). The outcome is achieved by 3D voxelized wireframes of multiple architectural forms with different styles and strengths through learning continuous latent distributions of VAE. Based on this work, Miguel et al. further tested parametric augmentation for a larger dataset of 3D geometries (Miguel et al., 2020). By contrast, some work concentrates on post-integration assisting in training according to the designer’s preference after the VAE generation in the whole production. For instance, Kahraman et al. produced a series of chairs with 3D voxelization while comparing multi-object VAE and GAN (Kahraman et al., 2021). This application discussed pre-defined criteria ranging from leisure to work scenarios.
Although VAEs are highly capable of incorporating 3D features, their current application is limited to generating specific targets with similar features, such as single buildings. They also produce low-resolution outputs because they use a variational loss to approximate the probability density by optimizing a lower bound on the log-likelihood function. Additionally, the high technical threshold hinders research and applications for non-experts in AI.
3.4. Approaches Using 3D-Aware Image Synthesis
@@
Most 3D-aware image synthesis methods utilize GAN-based models to sample latent vectors and decode them into target 3D representations. Thus, they offer 3D view-consistent rendering, efficient representation, and interactive editability, as well as integration capabilities with other deep learning generation techniques (Shi et al., 2023b), which could contribute to virtual architectural generation.
3.4.1. 3D-Aware Image Synthesis and Its Editability
@@
3D-aware synthesis shows promise for virtual architecture regarding representation conversion, controllability, and multi-modality with linguistic descriptions. As OrEl et al. demonstrated in StyleSDF, converting implicit representations to explicit geometry enables designers to comprehend and edit architectural elements with precision visually (OrEl et al., 2022). Similarly, Gu et al. showcased in StyleNeRF how explicit control over style attributes allows for creating diverse architectural designs with semantic editing capabilities (Gu et al., 2021). Recent advancements include Meng et al.’s introduction of a Colab configuration for creating buildings based on DreamField, which supports conditional text input, parameter adjustment, and style attributes  [^17] Source: https://github.com/shengyu-meng/dreamfields-3D. Additionally, Guida et al. have generated buildings with 3D representation by leveraging NeRF from generative images produced by Stable Diffusion (Guida et al., 2023). This method expands the potential for generating a wider variety of complex forms. This implementation underscores the practical application of 3D-aware synthesis in architectural design, providing designers with intuitive tools to realize their creative vision. However, concerns arise regarding its effectiveness, which is often limited to simple objects lacking high resolution and precise internal structures due to non-design datasets. To address these limitations, Guida et al. utilized integrated stable diffusion with text-to-image synthesis to supplement non-design datasets, effectively enhancing the diversity of form-finding (Guida et al., 2023). Nevertheless, the efficiency of generating complex architectural structures with both internal and external details remains constrained by the characteristics of deep generative models.
3.5. Emerging Approaches Using Diffusion Models
@@
3D diffusion models are capable of high quality with fine details in a robust and controlled way for 3D generation, particularly complex shapes with intricate details and specific attributes. Thus, their controllability and descriptive input will contribute to deep-learning-assisted virtual architectural generation in the future.
3.5.1. 3D Diffusion and Its Manipulability
Recently, 3D diffusion has gained a growing interest in generating 3D shapes due to its outperformance abovementioned. Despite the flexibility of conditional diffusion sampling, previous diffusion models are limited to sampling pixels of images. In contrast, Ben et al. have innovatively applied diffusion models to generate 3D models directly through applying denoising in high-quality 3D-aware image synthesis named DreamFusion (Poole et al., 2022). This work sparkes various follow-up approaches regarding text-to-3D (e.g., Magic3D (Lin et al., 2023), ProlificDreamer (Wang et al., 2023)) and image-to-3D (e.g., PVD (Zhou et al., 2021), Magic123 (Qian et al., 2023), MVDream (Shi et al., 2024)) with improved generation efficiency, resolution of geometry and texture. For instance, LION concentrates on greater flexibility regarding operation and application by leveraging conditional synthesis and shape interpolation within a combination of diffusion models and VAE (Zeng et al., 2022). While Magic123 with image input focuses on reconstructing a high-resolution and textured 3D mesh using joint 2D and 3D priors (Qian et al., 2023).
@@
Diffusion models facilitate the generation of 3D shapes with outperformed controllability in two ways: modifying style attributes and manipulating conditioned descriptive text. Firstly, most methods allow modification of specific style attributes, such as shape and texture, during the generation process (Gu et al., 2021; Niemeyer and Geiger, 2021; Jain et al., 2022). This progress suggests that diffusion models offer a robust and controlled approach to 3D shape generation, particularly for complex shapes with intricate details and specific attributes. Secondly, some burgeoning methods feature text-to-3D generation with human interpretative intention by manipulating text prompts (Chen et al., 2019; Poole et al., 2022; Liu et al., 2022; Lin et al., 2023; Nichol et al., 2022b). Earlier work utilized CLIP for 3D generation and manipulated text prompts, such as PointE (Nichol et al., 2022b). Comparably, current works have improved output results by incorporating stronger synthesis priors, such as DreamFusion (Poole et al., 2022).
3.5.2. 3D Diffusion Model Utilized in Architecture
Pioneering design works utilizing diffusion models have proliferated recently. While most of these works only utilize diffusion models to generate 2D images coupled with additional computational methods for 3D transposition, such as LoRA and NeRF (Guida et al., 2023). For instance, AIG studio’s work  [^18] Source: https://www.bilibili.com/video/BV1Qb411Z7UP/ has achieved a modeling framework through stable diffusion, generating plan drawings by relying on heatmaps of urban morphology and additional LoRA models. Comparably, an alternative method employs diffusion models for nonlinear shape interpolation to create a variety of 3D forms (Koh, 2023; Liu et al., 2024; Sebestyen et al., 2023). For example, one investigation utilized a 3D point cloud probabilistic diffusion model to generate Taihu stone in specified target shapes (Liu et al., 2024). Similarly, another study adopted volumetric density grids as the 3D representation within denoising diffusion models, providing a range of possibilities for architectural design (Sebestyen et al., 2023).
Some initial research in design has also leveraged diffusion models regarding integrating with human interpretation, interaction usability, and multimodality. An earlier work by Zhang developed a framework capable of compiling analytic semantic information from input texts, then generating 3D architectural form according to descriptive language (Zhang, 2020). In this framework, different usages of spaces were trained with adjacent matrices to understand the linguistic instructions. The approaches of modifying attributions and conditioned descriptive text broaden the usage case for other layman users by interpreting textual prompts. For instance, Magic 3D has developed a toolkit that offers advanced control over image instances of 3D generation regarding stylization and content based on text prompts (Lin et al., 2023).
Additionally, recent work has contributed to integrating tools to improve common tedious workflows (Guida, 2023), such as constantly transforming models between multiple software or methods. In this work, Guida developed a plug-in with a user interface embedded in Rhino, integrating multiple approaches of 3D form generation for designers (Guida, 2023). This work underscores the significance of multimodality for designer demands. The cutting-edge diffusion model’s high resolution, denoising, and rapid generation democratizes 3D geometry generation, providing access for individuals with varying levels of expertise to produce creatively.
4. Characteristics of Generated Virtual Architecture in Designer-AI Collaboration
@@ 4.2. Scope
The degree of scope depends on two aspects: the involvement of process scope in the design and deep learning-assisted design or human-assisted design. Firstly, involvement in more design phases represents a wider scope, including data training and processing, solution set generation, and evaluation (as illustrated in Section 2.3). For instance, Mueller conducted the entire scope of design work involving pre-processing datasets, refining and fine-tuning models, and evaluation based on GANs to test the impact of deep learning on design output (Mueller, 2023). By comparison, Zhang et al. merely applied GANs to generate a set of section images for 3D transposition without involving other phases, representing a specific scope (Zhang, 2019).
Secondly, the work involving human-assisted design has a wider scope than deep learning-assisted design regarding design generation since it accommodates additional processes beyond design generation, such as dataset augmentation (Miguel et al., 2019) and semantic labeling (Zhang, 2020). For instance, Huang et al. extended architects’ capabilities by combining the GAN generation with analytical interpretation processes of NLP (natural language processing) instead of solely focusing on design generation (Huang et al., 2021). Such research addresses the limitations of DL-assisted design generation and compensates for them by incorporating a wider scope of design perspectives. In contrast, most studies relying on generative images only investigate DL-assisted design without addressing other concerns (Zhang, 2019; Yu, 2020).
@@ 4.3. Operability
Operability refers to the extent of technical or design operation involved in design collaboration with DGMs. Low operability indicates the straightforward application of DGMs to design tasks, while high operability signifies the ability to operate within the technical space in a work with improved functionality and efficiency. Technical operations encompass various types across different models, including fine-tuning model parameters (Mueller, 2023), refining training procedures (Çakmak, 2022; Mueller, 2023; Sebestyen et al., 2021), customizing architecture (Kahraman et al., 2021), and adapting methodology applicability (Huang et al., 2023).
The degree of operability is influenced by the complexity of technical operations and the applicability of methodology. Firstly, the different focuses of design work contribute to varying levels of technical operations and the applicability of the methodology. For instance, Veselý demonstrated high operability by unifying methodology applicability and fine-tuning GAN models for different scales and scopes, including adjusting hyperparameters and optimizing the training process to achieve better convergence and output quality (Ondrej Vesel‘y, 2022). Similarly, Huang et al. explored the latent space potential by connecting structured segmentation of text and visual outputs in generative images of the diffusion model, establishing robust analytical frameworks operation (Huang et al., 2023). Secondly, different designs exhibit varying levels of applicability due to model divergence. Generally, VAEs require relatively high operability from professionals since it processes low-dimensional 3D feature data of buildings. In contrast, diffusion and GAN offer various levels of operability with flexibility ranging from application to fine-tuning, while 3D-aware synthesis requires an intermediate level of operability.
@@ 4.4. Autonomy
Autonomy refers to the level of automation in the generation process facilitated by deep learning techniques. High autonomy indicates a largely automated process driven by a unified workflow, while low autonomy involves manual intervention, requiring the connection of workflows and greater ease of use. In architectural design, most works exhibit relatively low autonomy to ensure quality and design accuracy under the guidance of designers. For example, Mueller validated the methodology across diverse scenarios by selectively retrieving spatial data and comparing the performance of proposed generative models (Mueller, 2023). However, recent advancements show a proliferation of high autonomy, especially in 3D generation, focusing on task efficiency (Kim et al., 2020; Lab, 2023). Nevertheless, there is a growing trend towards higher autonomy in design work to broaden access and enhance usability. For instance, Guida developed an integrated tool within the visual scripts of Grasshopper with high autonomy, combining all processes and modules in architectural design, providing convenient input and output modules across different approaches (Guida, 2023).
@@ 5. Research Agendas
In the previous sections, we explored the urgent need for HCI and DL-assisted methods in generated virtual architecture. We thoroughly examine the literature and elucidate four key focuses, as mentioned in Section 2.3. These key focuses serve as the foundation for our research agendas, formulated from a perspective of high-level questions (Fig. 9). Our investigation of current literature and works also reflects on these questions. Therefore, we advocate the following research agendas to promote research in this area.
Figure 9. A summary diagram illustrating the connections among key sections, including the overview (Section 2), generative approaches (Section 3), and characteristics of these approaches (Section 4). Accordingly, these high-level issues, identified within the four key sections of our survey, serve as the basis for further research agendas (Section 5).
@@ 5.1. Agency Between Designer and Machine in Data Processing and Training
Agency between humans and machines emerges as a significant feature in AI-assisted design and designer-AI collaboration. On the one hand, a human-driven agency faces challenges due to low-resolution and incomplete datasets, requiring human intervention (Shi et al., 2023a). On the other hand, an AI-driven agency raises concerns about intellectual property issues but offering higher efficiency (Guida, 2023). Our analysis of the survey identifies two main barriers to achieving a balance between human and AI agencies: flexible datasets and human-assisted machine learning.
@@ 5.2. Augmenting Communication Between Design and Artificial Intelligence
To benefit the design project’s fitting purpose, HCI researchers and designers should augment communication with each other by promoting tools or enhancing workflows between design applications and deep learning. We found two aspects closely related to this agenda: explainability in designer-AI collaboration and editability in the generative model.
Explainability. Explainability is one recognized principle of communication within designer-AI collaboration, essential for fostering trust in AI systems and guiding the design process and outcomes (Mohseni et al., 2021). However, a significant disparity exists in the developmental progress between HCI and the design field. For instance, HCI researchers have made tremendous strides in responsible AI for explainability, focusing on user experience, trustworthiness, and collaboration (Saeed and Omlin, 2023). The design field lags due to existing challenges and a lack of expertise among designers in negotiating input, process, and controlling output. According to our analysis of the survey, no one has addressed this critical issue. Regardless, recent studies (N=2) have primarily investigated methods for controlling output by manipulating detail in algorithms within generative models to achieve greater controllability in design outputs. Therefore, future research in HCI and design should delve deeper into explainability in design and generation processes building upon design principles.
Editability. Editability is the other key to communication, relying on timely feedback and adjustments to model 3D objects. Similarly, there is a big gap in development progress between the generation technique and design fields. Prior work has demonstrated editing autonomy in 3D shape generation techniques available to users. For example, Liu et al. proposed a user-centric 3D shape generative method by drawing a target model as 3D voxel grids (Liu et al., 2017). While a few works demonstrate some indirect approaches to editing 3D models with implicit representations (Hao et al., 2020; Ibing et al., 2021; Deng et al., 2021b; Zheng et al., 2021). According to our survey, although none of the architectural works mainly contribute to editing in 3D generation, one indicated a potential for variability in 3D form. This work generates infinite continuous forms by interpolating 3D features among multiple wireframe representations with compact implicit functions (Miguel et al., 2019). Through this interpolation with implicit representations, we could envision a broad use for virtual architecture in supporting itself with real-time dynamic morphology. Thus, the editability potential of implicit representation incorporating building form launches the future agenda for deep learning virtual architecture design.